# Flexing RISC-V Instruction Subset Processors (RISPs) to Extreme Edge

### v2

@@ Abstract
This paper presents a methodology for automatically generating processors that support a subset of the RISC-V instruction set for a new class of applications at Extreme Edge. The electronics used in extreme edge applications must be power-efficient, but also provide additional qualities, such as low cost, conformability, comfort and sustainability. Flexible electronics, rather than silicon-based electronics, will be capable of meeting these qualities. For this purpose, we propose a methodology to generate RISPs (RISC-V instruction subset processors) customised to extreme edge applications and to implement them as flexible integrated circuits (FlexICs). The methodology is unique in the sense that verification is an integral part of design. The RISP methodology treats each instruction in the ISA as a discrete, fully functional, pre-verified hardware block. It automatically builds a custom processor by stitching together the hardware blocks of the instructions required by an application or a set of applications in a specific domain. This approach significantly reduces the processor verification and its time-to-market. We generate RISPs using this methodology for three extreme edge applications, and embedded applications from the Embench benchmark suite, synthesize them as FlexICs, and compare their power, performance and area to the baselines. Our results show that RISPs generated using this methodology achieve, on average, 30% reductions in power and area compared to a RISC-V processor supporting the full instruction set when synthesized, and are nearly 30 times more energy efficient with respect to Serv - the world’s smallest 32-bit RISC-V processor. In addition, the full physical implementation of RISPs show up to 21% and 26% less area and power than Serv.
@@ 1 Introduction
There is a new class of emerging applications that require embedded electronics capable of sensing, computing, and communication providing qualitative features such as low cost, conformability, comfort, biocompatibility, and sustainability in addition to low power consumption. These applications are now described as “Extreme Edge” that extends beyond the edge computing today [12].
Extreme edge applications can be classified into two categories: 1) Short-lived extreme edge applications have short lifetimes. Item-level intelligence, Fast Moving Consumer Goods (FMCG) and single-use healthcare products are good examples of short-lived applications. Because electronics is part of the product itself, it will have a short lifetime, and any firmware/software update to the electronics will not be very likely. For instance, a smart dressing product has built-in electronics to monitor wound healing conditions. The lifetime of the dressing is in the range of days before it is replaced. The dressing and its built-in electronics will be disposed. 2) Long-lasting extreme edge applications have much longer lifespan. Smart textile, wearable healthcare, environmental/agriculture monitoring are good examples of long-lasting extreme edge applications. For instance, the electronics embedded into a smart garment can be monitoring the health status and performance of the person wearing it. The garment will be re-washed and re-used in a lifespan of many months before it wears out, and its built-in electronics may likely receive software/firmware updates during its lifetime. Table 1 summarizes the application domains and features of these two categories.
Conventional Silicon-based electronics struggle to meet the requirements of extreme edge because it does not satisfy many of the required features. For example, smart labels in asset tracking and smart packages in FMCG are ultra cost-sensitive applications where electronics embedded into them must not cost more than a few cents per unit to ensure economic viability. An ECG patch requires a comfortable and conformable form factor in electronics for the patient where conventional bulky electronics will cause discomfort. An implantable or ingestible device must contain electronics that are biocompatible and conformable. The volume of the extreme edge products in terms of units is enormous, measured in a few trillions annually [32, 35, 15, 14, 34, 18, 17, 33]. If electronic devices are embedded into such volume of products, they must be manufactured in sustainable semiconductor fabs that have a low environmental footprint.
Flexible electronics [23] is an alternative technology that meet many of these characteristics required by extreme edge. It aims to produce electronic components including sensors, compute, communication, and battery using low-cost printing and lithography techniques on flexible substrates (e.g., plastic or paper). They have an ultra-thin and flexible form factor and can be manufactured in sustainable fabs that use less water and energy, which improves the environmental footprint. For example, Pragmatic fabricates ultra low-cost flexible integrated circuits (FlexICs) using metal-oxide thin-film transistor technology on a polyimide wafer [31], as shown in Figure 1. The chips are ultra-thin, conformable, bendable and ultra-low cost [28], and the fabs that manufacture FlexICs have significantly lower carbon footprint per FlexIC than an equivalent Si chip [1]. Because the fabrication process is low-cost and fast (weeks rather than months) in manufacturing FlexICs, the development of custom FlexICs specifically designed for an application or domain can be faster and lower in cost than the Silicon technology.
@@
Short-lived Category Long-lasting Category
@@
The performance requirements of the two extreme edge categories are not intensive where a sensor or an array of sensors are connected to an analogue frontend to pre-process raw sensor data that are converted into digital data, which is processed by a digital backend to translate them into knowledge. The knowledge is stored and/or directly transmitted to the external world through a communication interface (e.g., RF). Data sampling rates required by sensors are 200 Hz[6], which implies that digital backend processing speeds can be in the range of Hz for many extreme edge applications or kHz for applications that require faster response.
Natively flexible smart systems [11] consisting of flexible electronic components, such as sensors, displays and batteries coupled with compute, memory and communication circuits, are a natural fit for extreme edge. Both general-purpose processors and domain-specific processing engines have been developed as FlexICs. For instance, the studies in [22, 21, 7, 3, 28, 5] demonstrated natively flexible 4-bit, 8-bit and 32-bit general-purpose processors whilst others proposed natively flexible domain-specific processing engines in the domains of RFID communication [20, 43], machine learning [24, 27, 25, 26, 16, 30], security [6], and healthcare [29, 46]. These natively flexible processors have been shown to operate at the kHz range reaching up to 100 kHz, which is sufficient to meet the performance requirements of extreme edge applications.
@@
Figure 1: FlexIC technology. (a) A 200 mm wafer on a polyimide substrate, (b) A FlexIC is an integrated circuit manufactured on a polyimide wafer substrate. The semiconductor material is a metal-oxide material called Indium-Gallium-Zinc-Oxide (IGZO) that can be manufactured at low temperatures. The image of the FlexIC diced from the polyimide wafer is shown. It is an ultra-thin (30  µm) and physically flexible or bendable. (Source: Pragmatic Semiconductor)
@@
Although designing a domain-specific processing engine is more efficient than a general-purpose processor, it takes a considerable time to design, test and verify the design if the ASIC approach is taken for domain-specific processing. An alternative is to use an application-specific instruction processor (ASIP) [41, 2], where the instruction sets are customized to the application or domain. However, this approach needs to generate not only the ASIP but also the software toolset, such as a compiler, a linker and a simulator associated with it.
@@
In this paper, we propose a methodology for automatically generating domain-specific processors for extreme edge applications. The proposed methodology relies on the open standard RISC-V [42] instruction set architecture (ISA) and generates processors that use only a subset of the RISC-V full ISA (e.g., RV32I). We observe (see Section IV) that extreme edge and extreme-edge like embedded applications only use 31-84% of the full RISC-V ISA. Thus, a processor supporting a subset of the full RISC-V ISA can be more efficient in terms of performance, power and area (PPA) because the logic related to the unused instructions is avoided. We call them RISC-V instruction subset processors or RISPs.
The RISP generation methodology takes a very different approach to a processor microarchitecture where verification is an integral part of design. The philosophy is to design a RISP from pre-verified building blocks rather than the traditional design-followed-by-verification approach. First, the semantics of all instructions in the ISA are written in RTL as instruction hardware blocks. Then, they are formally verified, and stored into a library. A RISP is formed by extracting the required instruction hardware blocks from the library, composing them in a modular manner and reducing the verification effort significantly. Since extreme edge applications do not require high-speed operating clock frequencies, single-cycle unpipelined RISPs suffice to meet their performance requirements.
@@
Figure 2: RISP generation methodology. STEP 1: Generation of domain-specific instruction subset from an application or a set of applications from a domain; STEP 2: Pulling instruction hardware blocks for the domain-specific instruction set from the pre-verified full ISA hardware library, and formation of ModularEX; STEP 3: Construction of the RISP by stitching ModularEX with fixed units like memory interfaces and the register file.
@@
The RISP generation methodology avoids the significant effort of generating proprietary software toolsets specific to ASIPs by relying on the RISC-V software eco-system. For short-lived extreme edge applications, we predict that RISPs do not need software updates due to the short life of electronics embedded in the extreme edge device. For long-lasting extreme edge applications, occasional software updates to RISPs are needed. This implies the updated application(s) need to be re-compiled but the generated code must be constrained to the instruction subset supported by the RISP. This requires the modification in the RISC-V compiler (e.g., gcc) backend to constrain code generation to the instruction subset of the RISP. This is a far more complex task than it sounds. Instead, we develop a tool that takes the code of the updated application(s) re-compiled to the full ISA, and re-writes the new instructions that are not in the instruction subset of the RISP in terms of the instructions in the instruction subset using Generative AI, more specifically, ChatGPT RISC-V Assembly plugin.
@@
We design and implement RISPs as FlexICs using Pragmatic’s 0.6µm FlexIC process. The proposed methodology can automatically generate RISPs that, on average, consume approximately 26% less power and occupy around 30% less area than a 32-bit RISC-V processor supporting the full ISA when synthesized. Compared to Serv, RISPs are on average 30 times more energy efficient.
@@
    Automatic generation of RISC-V instruction subset processors (RISPs) that implement only the subset of the RISC-V ISA targeting extreme edge applications, whilst still relying on the existing software eco-system
@@
    Design of a RISP using a pre-verified instruction hardware block library to reduce the verification time and time-to-market of RISPs significantly
@@
    A Generative AI based tool to re-target the new instructions in the updated application(s) to the instructions in the instruction subset supported by RISP for long-lasting extreme edge applications
The remainder of this paper is structured as follows: Section 2 discusses the related work in processor customization. Section 3 presents the proposed methodology of generating RISPs. Section 4 describes the experimental evaluation and also presents the physical implementation results of RISPs as FlexICs. Section 5 introduces the Generative AI-based code recompilation/re-target framework. Finally, Section 6 concludes the paper.
@@ 2 Related Work
There are two main approaches in the design of domain-specific processors: i) Top-down or subtractive customization approach where the customization is performed over an already designed processor by shaving off the gates that are not used by the target application; ii) Bottom-up or additive approach where the customization is performed at the design level by putting in the hardware blocks required by the applications.
@@
PDAT [4] and Bespoke processors design [10] are two examples of a subtractive customization approach. The main disadvantage of the subtractive approach is that the process of finding and removing a large number of unused gates and hardware blocks in a well-designed processor is limited because the processor is already designed in the most optimal way. For example, an arithmetic logic unit (ALU) is shared among many instructions and these subtractive approaches are conservative to remove the portions of the ALU even if some of the ALU-using instructions can be removed.
@@
ASIPs [13, 19, 41, 2] use an additive customization approach where they create the custom instruction set for the application or the domain, and then generate the ASIP executing the custom instruction set. A complete software toolchain (e.g., a compiler, assembler, linker and simulator) is also generated to support the ASIP, which is an overhead. They adopt conventional processor microarchitecture techniques, and verification is not integral part of the design. Conventional verification strategies must be used after generating ASIPs.
@@
On the other hand, verification is an integral part of the RISP generation methodology, and RISPs are created in a bottom-up approach using the novel pre-verified instruction hardware block library concept reducing the verification effort of the generated custom processor. Additionaly, RISPs rely on open-source software toolchain of the RISC-V software eco-system rather than producing a new software toolchain for a proprietary ISA like in ASIPs.
@@
Table 2: A high-level overview of instruction hardware blocks in the library.
@@
3 Methodology of RISP Generation
There are various ISAs some of which are proprietary (e.g., Arm, x86), while others, such as RISC-V, are open-source, modular and extendable [42]. Undoubtedly, RISC-V is one of the most popular ISAs that is rapidly established in various embedded domains. A significant design principle of the RISC-V ISA is its modularity by using instruction groups allowing computer architects to design custom processors based on the requirements of the target application. Additionally, the built-in support for custom instructions offers a significant advantage compared to other ISAs. Specifically, the ability to design custom instructions at both software and hardware levels for the acceleration of domain-specific applications ranks RISC-V as one of the most flexible open-source ISA for both academia and industry. The significant advancements in the acceleration of domain-specific applications, such as machine learning, signal processing, computer vision, cryptography, or graph analysis, have highlighted the necessity for an open-source, modular and extendable ISA, such as RISC-V.
While a modular ISA allows independent instruction groups to be separated and combined based on the application requirements, the modularity of the ISA implementation depends on how these instruction groups or extensions are designed in hardware. It is important to note that a modular ISA does not necessarily infer that the underlying implementation of the ISA is modular. Unless the ISA implementation inherently supports “instruction-level modularity”, the only way to remove a subset of the ISA is to modify the actual RTL design. However, removing support for a subset of the ISA from the actual ISA implementation is not straightforward and requires a deep understanding of the RTL implementation of the processor supporting the ISA. This process may introduce errors in the design and extensive verification is necessary to ensure that the modifications maintain the initial functionality. One of the key benefits of this work is to enhance the instruction-level modularity of the ISA implementation by defining instructions as independent, fully functional, pre-verified RTL blocks.

### v3

# Flexing RISC-V Instruction Subset Processors to Extreme Edge

@@ Abstract
This paper presents an automated approach for designing processors that support a subset of the RISC-V instruction set architecture (ISA) for a new class of applications at Extreme Edge. The electronics used in extreme edge applications must be area and power-efficient, but also provide additional qualities, such as low cost, conformability, comfort and sustainability. Flexible electronics, rather than silicon-based electronics, will be able to meet the above qualities. For this purpose, we propose a methodology for generating RISC-V instruction subset processors (RISSPs) tailored to these applications and implementing them as flexible integrated circuits (FlexICs). The methodology makes verification an integral part of the processor design by treating each instruction in the ISA as a discrete, fully functional, pre-verified hardware block. It automatically builds a custom processor by stitching together the instruction hardware blocks required by an application or a set of applications in a specific domain. We generate RISSPs using the proposed methodology for three extreme edge applications, and embedded applications from the Embench benchmark suite. When synthesized, RISSPs can achieve 8-to-43% reduction in area and 3-to-30% reduction in power compared to a processor supporting the full RISC-V ISA, and are also on average ~40 times more energy efficient than Serv - the world’s smallest 32-bit RISC-V processor. When physically implemented as FlexICs, the three extreme edge RISSPs achieve up to 42% area and 21% power savings with respect to the full RISC-V processor.
@@ 1 Introduction
A new class of emerging applications requires embedded electronics capable of sensing, computing, and communicating while also offering qualitative features such as low cost, conformability, comfort, biocompatibility, and sustainability in addition to low power consumption. This new class of applications is now described as “Extreme Edge” which extends beyond edge computing today [12].
Extreme edge applications can be classified into two main categories: 1) Short-lived applications that have short lifetimes. Item-level intelligence, Fast Moving Consumer Goods (FMCG) and single-use healthcare products are examples of short-lived applications. Since electronics are part of the product itself, they will have a short lifetime, and any firmware/software update to the electronics will not be very likely. For instance, a smart dressing product has built-in electronics to monitor wound healing conditions. The lifetime of the dressing is in the range of days before it is replaced. The dressing and its built-in electronics will be disposed. 2) Long-lasting applications that have much longer lifespan. Smart textile, wearable healthcare and environmental/agriculture monitoring are examples of long-lasting extreme edge applications. For instance, the electronics embedded into a smart garment can be monitoring the health status and performance of a person. The garment will be rewashed and reused in a lifespan of several months before it wears out, and its built-in electronics may receive software/firmware updates during its lifetime. Table 1 summarizes the application domains and features of these two categories.
@@
Category Short-lived Long-lasting
@@
Conventional silicon-based electronics are unable to meet the requirements of extreme edge applications because they do not satisfy many of the required features. For example, smart labels in asset tracking and smart packages in FMCG are ultra cost-sensitive applications where electronics embedded into them must not cost more than a few cents per unit to ensure economic viability. An electrocardiogram (ECG) patch requires a comfortable and conformable form factor in electronics for the patient where conventional bulky electronics will cause discomfort. An implantable or ingestible device must contain electronics that are biocompatible and conformable. The volume of extreme edge products in terms of units is enormous, measured in a few trillions annually [35, 36, 19, 21]. If electronic devices are embedded into such volume of products, they must be manufactured in sustainable semiconductor fabs that have a low environmental footprint.
@@
Figure 1: FlexIC technology. (a) A 200 mm wafer on a polyimide substrate, (b) A FlexIC is an integrated circuit manufactured on a polyimide wafer substrate. The semiconductor material is metal-oxide called Indium-Gallium-Zinc-Oxide (IGZO) that can be manufactured at low temperatures. The image of the FlexIC diced from the polyimide wafer is shown. It is an ultra-thin (30  µm) and physically flexible or bendable.
Flexible electronics [26] is an alternative technology that meets many of the qualities required by extreme edge. The aim is to produce electronic components that include sensors, compute engines, communication interfaces, batteries using low-cost printing, and lithography techniques on flexible substrates (e.g., plastic or paper). For example, flexible integrated circuits (FlexICs) [34] can be fabricated using metal-oxide thin-film transistor technology on a polyimide wafer, as shown in Figure 1.
The chips are ultra-thin, conformable, bendable and ultra-low cost [31], and the semiconductor fabs that manufacture FlexICs have significantly lower carbon footprint per FlexIC than an equivalent Si chip [1]. Because the fabrication process is low-cost and fast, the development of custom FlexICs specifically designed for an application or domain can be faster and lower in cost than Silicon technology.
The performance requirements of the two extreme edge categories are not demanding where a sensor or an array of sensors are connected to an analogue frontend to pre-process sensor raw signals that are converted into digital data and handled by a digital backend to translate them into knowledge. The knowledge is stored and/or directly transmitted to the external world through a communication interface (e.g., RF). Data sampling rates required by sensors are 200 Hz[6], which implies that digital backend processing speeds can be in the range of Hz for many extreme edge applications or kHz for applications that require faster response.
Natively flexible smart systems [11] consisting of flexible electronic components, such as sensors, displays and batteries coupled with compute engines, memory and communication units, are a natural fit for extreme edge. Both general-purpose processors and domain-specific processing engines have been developed as FlexICs. For instance, the studies in [25, 24, 7, 3, 31, 5] demonstrated natively flexible 4-bit, 8-bit and 32-bit general-purpose processors whilst others proposed natively flexible domain-specific processing engines in the domains of RFID communication [23, 42], machine learning [27, 30, 28, 29, 20, 33], security [6], and healthcare [32, 43]. These natively flexible processors have been shown to operate at the kHz range reaching up to 100 kHz, which is sufficient to meet the performance requirements of extreme edge applications.
@@
Although designing a domain-specific processing engine is more efficient than a general-purpose processor, it takes a considerable time to design, test and verify the design if the ASIC approach is taken for domain-specific processing. An alternative is to use an application-specific instruction processor (ASIP) [40, 2], where the instruction sets are customized based on the application or domain. However, this approach needs to generate not only the ASIP but also the associated software toolset, such as a compiler, linker and simulator.
@@
In this paper, we propose a methodology for automatically generating domain-specific processors for extreme edge applications. The proposed methodology relies on the open standard RISC-V [41] instruction set architecture (ISA) and generates processors that use only a subset of the RISC-V full ISA (e.g., RV32I). We observe (see Section 4) that extreme edge and extreme edge-like embedded applications only use 24-86% of the full RISC-V ISA. Thus, a processor supporting a subset of the full RISC-V ISA can be more efficient in terms of performance, power and area (PPA) because the logic related to the unused instructions is avoided. We define them as RISC-V Instruction SubSet Processors or RISSPs.
@@
Figure 2: RISSP generation methodology. STEP 0: Development of the pre-verified full ISA hardware library, which is the one-time development effort of the methodology; STEP 1: Generation of domain-specific instruction subset from an application or a set of applications from a domain; STEP 2: Pulling instruction hardware blocks for the domain-specific instruction set from the pre-verified full ISA hardware library, and formation of ModularEX; STEP 3: Construction of the RISSP by stitching ModularEX with fixed units like memory interfaces and the register file
The RISSP generation methodology takes a very different approach to a processor microarchitecture where verification is an integral part of the design process. The main concept is to design a RISSP from pre-verified building blocks rather than the traditional design-followed-by-verification approach. First, the semantics of all instructions in the ISA are written in RTL as instruction hardware blocks. Then, they are formally verified, and stored into a library. A RISSP is formed by extracting the required instruction hardware blocks from the library, composing them in a modular manner and reducing the verification effort significantly. Since extreme edge applications do not require high-speed operating clock frequencies, single-cycle non-pipelined RISSPs suffice to meet their performance requirements.
@@
The RISSP generation methodology avoids the significant effort of generating proprietary software toolsets specific to ASIPs by relying on the RISC-V software ecosystem. For short-lived extreme edge applications, we predict that RISSPs do not need software updates due to the short life of electronics embedded in the extreme edge device. For long-lasting extreme edge applications, occasional software updates to RISSPs are needed. In this context, the updated application(s) need to be recompiled but the generated code must be constrained to the instruction subset supported by the target RISSP. This requires several modifications in the RISC-V compiler (e.g., gcc) backend to constrain code generation to the instruction subset of the RISSP. However, these modifications involve various challenges at multiple levels of the compiler stack. Instead, we develop a Generative AI-based tool that receives the assembly code of the updated application(s) recompiled to the full ISA, and rewrites the new instructions that are not supported by the RISSP in terms of the instruction subset.
@@
The proposed methodology can automatically generate RISSPs that are 8-to-43% smaller and 3-to-30% lower power, when synthesized, compared to a processor supporting the full RISC-V ISA, also generated by the same methodology, and on average about 40 times more energy efficient than Serv [38]. When physically implemented as FlexICs, the extreme edge RISSPs achieve up to 42% area and 21% power savings relative to the full RISC-V processor.
@@
    Automatic generation of RISC-V instruction subset processors (RISSPs) that implement only the subset of the RISC-V ISA targeting extreme edge applications, whilst still relying on the existing software ecosystem
@@
    Design of a RISSP using a pre-verified instruction hardware block library to significantly reduce the verification time and time-to-market of RISSPs
@@
    A Generative AI based tool to retarget the new instructions in the updated application(s) to the instructions in the instruction subset supported by RISSP for long-lasting extreme edge applications
The remainder of this paper is structured as follows: Section 2 discusses the related work in processor customization. Section 3 presents the proposed methodology of generating RISSPs. Section 4 describes the experimental evaluation and also presents the physical implementation results of RISSPs as FlexICs. Section 5 introduces the Generative AI-based code recompilation/retarget framework. Finally, Section 6 concludes the paper.
@@ 2 Related Work
There are two main approaches in the design of domain-specific processors: i) Top-down or subtractive customization approach where the customization is performed over an already designed processor by removing the gates that are not used by the target application; ii) Bottom-up or additive approach where the customization is performed at the design level by placing in the hardware blocks required by the applications.
@@
PDAT [4] and Bespoke processors [10] are two examples of a subtractive customization approach. PDAT converts a baseline processor into a standard netlist and uses property assertions to determine whether a specific gate is utilized during the execution of an embedded application. It exploits gate-level property checking to replace unused cells with wires if the relevant property is not asserted. Bespoke processors focus on the development of processing hardware from a baseline processor customized to a specific application. It uses input-independent gate activity analysis to remove the gates from the baseline processor that do not toggle during execution, resulting in a bespoke processor. The main disadvantage of the subtractive approach is that the opportunity of finding and removing a large number of unused gates and hardware blocks in a baseline processor is limited because processors are typically designed as lean as possible. For example, an arithmetic logic unit (ALU) is shared among many instructions, and these subtractive approaches are conservative to remove the portions of the ALU even if some of the ALU-using instructions can be removed.
@@
ASIPs [18, 22, 40, 2] use an additive customization approach in which they create the custom instruction set for the target application or domain and then generate the ASIP that executes the custom instruction set. A complete software toolchain (e.g., a compiler, assembler, linker and simulator) is also generated to support the ASIP, which is an overhead. They adopt conventional processor microarchitecture techniques, and verification is not an integral part of the design. Conventional verification strategies must be used after generating ASIPs.
@@
On the other hand, verification is an integral part of the RISSP generation methodology, and RISSPs are created in a bottom-up approach using the novel pre-verified instruction hardware block library concept reducing the verification effort of the generated custom processor. In addition, RISSPs rely on open-source software toolchain of the RISC-V software ecosystem rather than producing a new software toolchain for a proprietary ISA like in ASIPs.
3 Methodology of RISSP Generation
RISC-V [41] is one of the most popular ISAs that is rapidly established in various embedded domains. A key design principle of the RISC-V ISA is its modularity allowing computer architects to design custom processors based on the requirements of the target application. Additionally, the built-in support for custom instructions offers a significant advantage compared to other ISAs. While a modular ISA allows independent instruction groups to be separated and combined based on the application requirements, the modularity of the ISA implementation depends on how these instruction groups or extensions are designed in hardware. It is important to note that a modular ISA does not necessarily infer that the underlying implementation of the ISA is modular. Unless the ISA implementation inherently supports “instruction-level modularity”, the only way to remove a subset of the ISA is to modify the actual RTL design. However, removing support for a subset of the ISA from the actual ISA implementation is not straightforward and requires a deep understanding of the RTL implementation of the processor supporting the ISA. This process may introduce errors in the design and extensive verification is necessary to ensure that the modifications maintain the functionality.
One of the key benefits of this work is to enhance the instruction-level modularity of the ISA implementation by defining instructions as independent, fully functional, pre-verified RTL blocks.
@@ 3 Methodology of RISSP Generation
Table 2: A high-level overview of instruction hardware blocks of the RISC-V RV32I/E ISA